Circular No. 7881

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002ci IN UGC 10301 P. Martin and W. D. Li, University of California at Berkeley (UCB),

report the LOTOSS (cf. IAUC 7514) discovery of an apparent supernova on unfiltered CCD images taken with the 0.8-m Katzman Automatic Imaging Telescope (KAIT) on Apr. 19.5 (mag ∼ 17.2) and 20.4 UT (mag ∼ 17.1). SN 2002ci is located at α = 16^{h}15^{m}37^{s}.95, δ = +31^{o}19′17′′.4 (equinox 2000.0), which is 2′′.0 east and 0′′.1 north of the nucleus of UGC 10301. A KAIT image taken on Apr. 4.5 showed nothing at this position (limiting mag ∼ 19.0). R. Chornock, W. D. Li, and A. V. Filippenko, UCB, report that inspection of CCD spectra (range 330–1000 nm) obtained on Apr. 20 with the Shane 3-m reflector at Lick Observatory reveals that the object is a supernova of type Ia. The spectral-feature age (cf. IAUC 7774) is −1 ± 2 days relative to maximum light. Strong Na I D absorption at the redshift of the host galaxy (0.4-nm equivalent width) implies that this object may be somewhat reddened.

SUPERNOVA 2002cg IN UGC 10415 T. Matheson, S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian

Center for Astrophysics, report that a spectrum (range 370–750 nm) of SN 2002cg (cf. IAUC 7877), obtained by M. Calkins on Apr. 15.36 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a supernova of type I, but the exact classification is uncertain. It resembles SN 2002ap about a month past discovery. Chornock et al. remark that spectra (obtained at Lick Observatory, as above) indicate that SN 2002cg is a relatively normal supernova of type Ic.

SUPERNOVA 2002ch IN ANONYMOUS GALAXY Chornock et al. also note that spectra (obtained as above) show that

SN 2002ch (IAUC 7879) is a type-II supernova. Broad Hα emission greatly dominates over the absorption component. A redshift of ≈ 10 200 km/s is derived from narrow nebular emission lines associated with the host galaxy.

COMET C/2002 F1 (UTSUNOMIYA) Visual m_{1} estimates (B = binoculars): Apr. 9.13 UT, 6.6 (K. Hornoch,

Lelekovice, Czech Republic, 10×80 B); 11.38, 5.6 (J. Bortle, Stormville, NY, 10×50 B); 16.12, 4.1 (W. Hasubick, Buchloe, Germany, 10×50 B); 19.09, 4.5 (B. H. Granslo, Fjellhamar, Norway, 7×50 B).

2002 April 21  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green